OpenStudy (anonymous):

Could someone give me an example of the pythagoras theorem?? I dont really understand how to do it....

OpenStudy (hoblos):

 a^2 = b^2 + c^2

OpenStudy (anonymous):

a^2+b^2=c^2 ex. a=3 b=4 c=5 3^2+4^2=5^2

OpenStudy (anonymous):

c is the hypothenuse. 9+16=25

OpenStudy (anonymous):